OpenStudy (anonymous):

what is x2 - 14x + 49 = 0

OpenStudy (anonymous):

i believe it is called an "equation"

OpenStudy (anonymous):

please apply ax^2 +bx + c = 0

OpenStudy (anonymous):

SEE THE SOLUTION OF THIS QUESTION IS FOLLOWING

OpenStudy (anonymous):

a(x^2) + bx + c = 0 x is the only variable and a b and c are constants

OpenStudy (anonymous):

now

OpenStudy (anonymous):

-b pm

OpenStudy (anonymous):

pm means plus or minus both ok

OpenStudy (anonymous):

(-b + sqrt of(b^2-4ac))/2a

OpenStudy (anonymous):

x^2-14x+49=0 find two numbers that when you multiply you get 49 and when you add you get -14 those numbers are -7 and -7 x^2-7x-7x+49=0 x is a common factor x(x-7) -7(x-7) = 0 x-7=0 x=7